CH/DOC/94/12

Equipment at the ANU

Inventory after July ’94 Running

Author: R.P. Ward

Date: 21 Sep 1994

ANU Equipment Inventory, July 1994

Mounted Targets

Material Thickness [μg cm^{−2}] Number Remarks
12C  100  1
12C  200  2
12C  450  2
12C  300  4
12C  180  3
SiO_{2}  50  3
Flash gold  5 on 10  2 Very fragile
197Au  125  2

Unmounted Target Material

Material Thickness [μg cm^{−2}] Area [cm^{2}] Remarks
12C  10  60
SiO_{2}  50  72
Flash gold  5 on 10  24
197Au  125  6

Exabytes

32 × 112m unused. We also have a large number of spare copies, which I suggest we start recycling (or ship home for recycling, if people want to write data to new

tapes).

Anode Grids

7 × 100 μm (1 horizontal, 6 vertical). 2 × 20 μm (1 horizontal, 1 vertical). 5 × 20 μm (all horizontal).

Shield Grids

2 × frames. 3 × 20 μm. 2 × 100 μm.

Windows

8 × 3.5 μm (1 used).

Detectors

9 Hammamatsu detectors (numbers 4, 5, 6, 9, 10, 12, 13, 14 and 15. Number 15 has 3 strips missing; the others have no strips missing). 3 new and 10 used Micron

detectors.

Shipping

We have been shipping equipment to the ANU for temporary use via Local Port Practice at Manchester. However, when the return shipment came to Customs at Birmingham ,it was delayed for three weeks. I thus sugggest that we ensure that

outward and return shipping is always done via Manchester in future.